OpenStudy (anonymous):

Locate the saddle point of the function. f(x,y) = x^(2) − 6x + y − e^(y)

OpenStudy (anonymous):

Can you locate all critical points?

OpenStudy (anonymous):

You want \(f_x =0\) and \(f_y=0\)

OpenStudy (anonymous):

i get confused with e

OpenStudy (anonymous):

You can't differentiate it?

OpenStudy (anonymous):

\[ \frac{d}{dy}e^y=e^y \]

OpenStudy (anonymous):

so would fy = 1 - ye^(y) ?

OpenStudy (anonymous):

No, there shouldn't be a \(y\) in front.

OpenStudy (anonymous):

just 1 - e^(y) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would fyy = e^(y)

OpenStudy (anonymous):

where is the -?

OpenStudy (anonymous):

oops -e^(y)

OpenStudy (anonymous):

okay so right now i have: fx = 2x - 6 fy = 1 - e^(y) fxx = 2 fxy = 0 fyy = -e^(y)